OpenStudy (anonymous):

Solve the equation. 1.3x+ 2.8= 0.5x -4.48 The answer is -9.1 but how do I get it?

hero (hero):

1.3x+ 2.8= 0.5x -4.48 1.3x - 0.5x = -4.48 - 2.8 0.8x = -7.28 x = -7.28/0.8 x = -9.1

OpenStudy (mathhelp346):

1.3x+2.8= 0.5x -4.48 +4.48 +4.48 1.3x+7.28=0.5x -0.5x -0.5x .8x+7.28=0 -7.28 -7.28 8x=-7.28 then multiply both sides by 1/8